OpenStudy (anonymous):

What is the pH of a solution with a [H+] of 1.0 × 104 M? (Points : 1) 2.0 3.0 4.0 8.0

OpenStudy (anonymous):

A cup of black coffee has a pH of 5. What is the hydrogen ion concentration of this cup of coffee? (Points : 1) 3.0 × 10 negative5 M 1.0 × 10negative5 M 4.0 × 10 negative4 M 4.0 × 10 negative6 M

OpenStudy (lissiewalkin):

I think what you do is do punch in pH = -log(H+ concentration) so pH = -log(1.0 X 10^4) im assuming that 104 is 10^4?

OpenStudy (anonymous):

its 10^-4

OpenStudy (lissiewalkin):

I think the first one should be 4 if you put it in the calculator pH = -log(1.0 X 10^-4)

OpenStudy (anonymous):

is it D?

OpenStudy (lissiewalkin):

you didnt put letters for your options but i guess it would be c) 4.0

OpenStudy (anonymous):

thanks could you help with the other one

OpenStudy (lissiewalkin):

i believe it is 1.0 × 10negative5 M because the formula to find pH is -log(H+ concentration) so you plug the values in 5 = -log(H+) ---(H+ is what you are trying to find) then you get rid of the -log by doing 10^-5 and 10^-5 is 1.0 x 10 negative 5

OpenStudy (anonymous):

okay thank you